Meinongian Logics

Jacek Paśniczek

Alexius Meinong, an Austrian philosopher (1853-1920), is particularly known as an”inventor” of an abundent, all-encompassing ontology of objects understood as possible objects of thought. Among these objects even incomplete and impossible ones, like the round square, find their place. Meinong coined the slogan: There are objects of which it is true that they do not exist . Bertand Russell strongly critisised Meinong’s theory of objects accusing it of inconsistency. Since the critics seemed to be decisive, the theory was rejected by the majority of philosophers and was almost forgotten for a long time. Only recently there has been a renaissance in Meinong studies. In the last three decades several theories of objects which are formalisations of Meinong s views or which are inspired by them have been independently developed by Parsons, Castaneda, Routley, Rapaport, Zalta, Jacquette, and Paśniczek. These theories show that Meinong’s views can be interpreted in a consistent way and this interpretation leads to highly interesting formal ontology. Some of the theories may be considered as logics sensu stricto, just as Meinongian logics (Meinong himself did not create any logic in the proper meaning of the term). Generally, the Meinongian logic can be viewed as a logic of non-existent objects. However, as opposed to other kinds of logics allowing for nonexistent objects (e.g. free logics), this logic associates with nonexistent objects a non-trivial predication. According to Meinongian logic the round square is round and is a square but this object basically does not possess any other properties (this fact perfectly accords with our language intuitions).

The aim of this paper is to present main principles of the Meinongian logic, to show its philosophical applications, and to relate it to other types of logics.

Address: Maria Curie-Skłodowska University, Lublin, Poland

E-mail: jpasnicz@platon.umcs.lublin.pl